OpenStudy (anonymous):

find the complex zeros of the polynomial function. write f in factored form f(x)=x^4+12x^2+11

jigglypuff314 (jigglypuff314):

try thinking about it like how would you have factored x^2 + 12x + 11 ?

OpenStudy (anonymous):

do you put the like terms together

jigglypuff314 (jigglypuff314):

mmm I don't see how there are like terms?

OpenStudy (anonymous):

yeah i hate math so im confused

jigglypuff314 (jigglypuff314):

factor x^2 + 12x + 11 do you know how to do that?

OpenStudy (anonymous):

no

jigglypuff314 (jigglypuff314):

well how about like this x^2 + 6x + 5 = (x + 5)(x+1) because 5+1=6 (the middle term) and 5*1=5 (the third term)

OpenStudy (anonymous):

so it would be (x+11)(x+1)

jigglypuff314 (jigglypuff314):

yep! :) then x^4 + 12x^2 + 11 is the same idea just replace x with x^2 in the factored form :) like x^4 + 6x^2 + 5 = (x^2 + 5)(x^2 + 1)

OpenStudy (anonymous):

ok so x^4+12x^2+11=(x^4+12)(x^4+11)

jigglypuff314 (jigglypuff314):

mmm not quite :) x^4 + 6x^2 + 5 = (x^2 + 5)(x^2 + 1) ^ left with x^2 not x^4

OpenStudy (anonymous):

so its (x^2+12)(x^2+11) yeah now im confused why is it ^2 and not ^4

jigglypuff314 (jigglypuff314):

(x^2 + 11)(x^2+1) because when you FOIL it (distribute it out) you would get (x^2)(x^2) + 11x^2 + 1x^2 + (1)(11) which simplifies to x^4 + 12x^2 + 11

OpenStudy (anonymous):

okay no i get it we are using the FOIL method

OpenStudy (anonymous):

so in factored form would be (x^2)(x^2) + 11x^2 + 1x^2 + (1)(11)

jigglypuff314 (jigglypuff314):

mmm that would be the factored form distributed factored form would be (x^2+11)(x^2+1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so this is factored form using complex zeros (x^2+11)(x^2+1)

jigglypuff314 (jigglypuff314):

then to find the complex zeros find x for each when x^2+11=0 and x^2+1=0

OpenStudy (anonymous):

ok so f(x)=

jigglypuff314 (jigglypuff314):

http://www.wolframalpha.com/input/?i=0%3Dx%5E4%2B12x%5E2%2B11

OpenStudy (anonymous):

okay they gave me f(x)=(x^2+1)(x^2+11)

jigglypuff314 (jigglypuff314):

hmm?

OpenStudy (anonymous):

is it wrong

jigglypuff314 (jigglypuff314):

no?

OpenStudy (anonymous):

yes whats the question

jigglypuff314 (jigglypuff314):

write f in factored form f(x)=x^4+12x^2+11 would be f(x)=(x^2+1)(x^2+11) then the other half of the question find the complex zeros of the polynomial function. would be find x for each when x^2+11=0 and x^2+1=0

OpenStudy (anonymous):

then it says use the complex zeros to write f in factor form

jigglypuff314 (jigglypuff314):

mmm then yeah, factored form is f(x)=(x^2+1)(x^2+11)

OpenStudy (anonymous):

ok